A goldsmith combined pure gold that cost $400 per ounce with an alloy of gold that costs $150 per ounce. How many ounces of each were used to make 50 oz of gold alloy that that costs $250 per ounce?
Answer by rfer(10432)

400x+150(50-x)=250*50
400x+7500-150x+12500
250x=5000
x=20
$400=20 oz
$150=30 oz
------------
400*20=8000
150*30=4500
total= 12500
-------------
12500/50=250